# Neural Combinatorial Optimization for Real-World Routing

### v1

@@
Abstract.
Vehicle Routing Problems (VRPs) are a class of NP-hard problems ubiquitous in several real-world logistics scenarios that pose significant challenges for optimization. Neural Combinatorial Optimization (NCO) has emerged as a promising alternative to classical approaches, as it can learn fast heuristics to solve VRPs. However, most research works in NCO for VRPs focus on simplified settings, which do not account for asymmetric distances and travel durations that cannot be derived by simple Euclidean distances and unrealistic data distributions, hindering real-world deployment. This work introduces RRNCO (Real Routing NCO) to bridge the gap of NCO between synthetic and real-world VRPs in the critical aspects of both data and modeling. First, we introduce a new, openly available dataset with real-world data containing a diverse dataset of locations, distances, and duration matrices from 100 cities, considering realistic settings with actual routing distances and durations obtained from Open Source Routing Machine (OSRM). Second, we propose a novel approach that efficiently processes both node and edge features through contextual gating, enabling the construction of more informed node embedding, and we finally incorporate an Adaptation Attention Free Module (AAFM) with neural adaptive bias mechanisms that effectively integrates not only distance matrices but also angular relationships between nodes, allowing our model to capture rich structural information. RRNCO achieves state-of-the-art results in real-world VRPs among NCO methods. We make our dataset and code publicly available at https://github.com/ai4co/real-routing-nco.
Neural Combinatorial Optimization, Vehicle Routing Problem, Reinforcement Learning, Dataset
copyright: noneccs: Mathematics of computing Combinatorial optimizationccs: Computing methodologies Sequential decision makingccs: Applied computing Transportationccs: Applied computing Supply chain managementccs: Computing methodologies Reinforcement learning
1. Introduction
@@
Vehicle routing problems (VRPs) are combinatorial optimization (CO) problems that represent foundational challenges in logistics and supply chain management, directly impacting operations across diverse sectors, including last-mile delivery services, disaster response management, and urban mobility. These NP-hard optimization problems require determining optimal routes for a fleet of vehicles while satisfying various operational constraints. While several traditional methods have been developed over decades (Laporte and Nobert, 1987; Vidal, 2022; Wouda et al., 2024; Perron and Furnon, 2023; Applegate et al., 2003; Wouda and Lan, 2023), these often face significant challenges in real-world applications. Their computational complexity makes them impractical for large-scale and real-time applications. Moreover, they often require careful parameter tuning, problem-specific adaptations, significant domain expertise, and lengthy development. With the global logistics market exceeding the $10$ trillion USD mark in 2025 (Research and Markets, 2024), improvements in routing efficiency can yield substantial cost savings and environmental benefits.
@@
Neural Combinatorial Optimization (NCO) has emerged as a promising paradigm for solving CO problems such as the VRP (Bengio et al., 2021; Wu et al., 2024a). By automatically learning heuristics directly from data, i.e., by using Reinforcement Learning (RL), NCO approaches for VRPs can potentially overcome the limitations of traditional methods by providing efficient solutions without requiring extensive domain expertise and by providing more scalable solutions (Kool et al., 2019; Zhou et al., 2023). Recent advances in NCO have demonstrated impressive results on synthetic VRP instances, suggesting the potential for learning-based approaches to achieve significant impact in real-world logistics optimization (Kwon et al., 2020; Luo et al., 2024b; Ye et al., 2024b; Hottung et al., 2025b).
Figure 1. [Left] Most NCO works consider simplified Euclidean settings. [Right] Our work models real-world instances where durations and travel times can be asymmetric.
@@
However, the transition from synthetic to real-world VRP applications faces two critical challenges. Firstly, most existing NCO research primarily relies on simplified synthetic datasets for both training and testing that fail to capture the complexities of real-world routing scenarios, particularly asymmetric travel times and distances arising from road networks with diverse conditions (Osaba, 2020; Thyssens et al., 2023a). Hence, a comprehensive framework for real-world data generation is needed to bridge this gap. Secondly, most current NCO architectures are mostly based on the node-based transformer paradigm (Vaswani et al., 2017) and, as such, are not designed to effectively and efficiently embed the rich edge features and structural information present in real-world routing problems, limiting their practical applicability (Kwon et al., 2021). A new neural approach to encode information as durations and distances is thus needed to fill the “sim-to-real” gap from synthetic settings to the real-world.
Our Real Routing NCO (RRNCO) aims to bridge the NCO gap between synthetic settings and real-world – as shown in Fig. 1 – by addressing these fundamental gaps through two core contributions. First, on the data side, we introduce a comprehensive real-world VRP dataset for NCO comprising diverse locations and with distance and duration matrices from 100 cities, capturing essential real-world complexities, incorporating actual routing data from the OpenStreetMap Routing Engine (OpenStreetMap contributors, 2025). Second, on the modeling side, we propose a novel architecture enhancing attention-free networks through contextual gating and adaptive bias mechanisms, enabling more effective processing of both node and edge features while capturing crucial structural relationships between locations.
The key contributions of RRNCO are summarized as follows:
  • We introduce a new extensive VRP dataset that bridges the gap between synthetic and real-world routing scenarios for both training and testing, featuring diverse city topologies and actual routing information.
  • We propose a novel NCO architecture incorporating contextual gating and adaptive bias mechanisms that effectively processes both node and edge features while capturing structural relationships.
  • We perform a comprehensive empirical evaluation demonstrating state-of-the-art performance on real-world VRP instances among NCO methods.
  • We make our dataset and code implementation open-source to facilitate further research on NCO for real-world impact.
2. Related Works
Neural Combinatorial Optimization (NCO). Neural approaches to combinatorial optimization have emerged as a promising paradigm for solving complex problems without extensive domain expertise (Bengio et al., 2021). NCO approaches can be broadly categorized into two main paradigms: construction methods and improvement methods.
Construction methods build solutions sequentially by learning policies that select elements step by step. Early approaches include Vinyals et al. (2015), who introduces pointer networks for sequence-to-sequence learning on CO problems, later enhanced by Bello et al. (2016) through policy gradient methods for reinforcement learning. Constructive approaches developments including autoregressive approaches that update decisions based on partial solutions (Kool et al., 2019; Kwon et al., 2020; Kim et al., 2022), non-autoregressive methods predicting a edge-based heatmaps in one-shot (Joshi et al., 2020; Fu et al., 2021; Kool et al., 2022; Sun and Yang, 2024; Li et al., 2024a, 2025) and (autoregressive) population-based approaches (Hottung et al., 2025a; Grinsztajn et al., 2024; Chalumeau et al., 2024) that maintain solution diversity , with genetic algorithms also being used for this purpose (Kim et al., 2025).
Improvement/search methods, on the other hand, start with initial solutions and iteratively enhance them through learned local adjustments. Notable works include learning improvement operators (Hottung and Tierney, 2019), neural large neighborhood search (Ma et al., 2021), and hybrid approaches combining neural networks with traditional local search (Ma et al., 2024). These methods have been particularly successful in cases where longer computation times are acceptable, with test-time adaptation techniques (Hottung et al., 2021; Choo et al., 2022; Camerota Verdù et al., 2025; Son et al., 2023) allowing for solution refinement.
Autoregressive constructive approaches – similar to the modern large language models (LLMs) (Dubey et al., 2024; Team, 2024) – often provide high solution quality with fast speed and do not require significant domain expertise, and are thus the scope of our proposed model.
Vehicle Routing Problem (VRP) Datasets. Over decades of research, evaluation datasets have been introduced for standardizing the evaluation of VRPs across the operations research and NCO community, such as the traveling salesman problem (TSP) library TSPLIB (Reinelt, 1991), the capacitated VRP library CVRPLIB (Lima et al., 2014) and their related extensions (Solomon, 1987; Uchoa et al., 2017; Queiroga et al., 2022; Thyssens et al., 2023b). These works are, however, mostly for evaluation only and often only include Euclidean coordinates. While some works provide more realistic instances (Osaba, 2020; Lee and Chae, 2021), these are usually in the order of few hundred and are not enough for training deep learning models, which require large datasets with fast online data generation as the (Euclidean) ones in Nazari et al. (2018); Kool et al. (2019); Bi et al. (2022). More related works to our scope include larger-scale generators based on realistic routing data (Duan et al., 2020; Ali and Saleem, 2024). However, these works employ commercial API providers, do not include duration matrices, cannot be generated online during training (i.e., Ali and Saleem (2024) reports 20+ seconds per instance for sizes greater than 300 nodes), and most importantly, do not make their code and datasets available to the research community. Our RRNCO aims to fill these gaps by providing a fully open-licensed and local API with both distance and duration data, fast online sampling for deep RL, and open availability to the research community.
NCO for VRPs. The application of neural approaches to VRPs has evolved significantly, pioneered by Nazari et al. (2018) introduce embeddings dynamic elements of VRP instances to pointer networks, enabling the handling of varying customer demands and vehicle capacities. Most recent works are based on the transformer-based architecture firstly proposed from Kool et al. (2019) – either in encoder-decoder (Kwon et al., 2020; Luo et al., 2024a) or decoder-only fashion (Drakulic et al., 2024b) – that better capture problem structure through attention mechanisms. NCO for VRPs has also expanded to several problems beyond classic TSP and CVRP (Bogyrbayeva et al., 2023; Bi et al., 2024; Berto et al., 2024a; Luttmann and Xie, 2025; Chen et al., 2024; Li et al., 2024b; Xiao et al., 2024).
Recent research has pursued several directions to enhance solution quality and generalization. Learning with problem re-encoding (Bdeir et al., 2022; Drakulic et al., 2024b; Luo et al., 2024a) enables better adaptation to partial solutions during construction. Several autoregressive approaches have been proposed, including population-based methods (Grinsztajn et al., 2024), mixtures of experts (Zhou et al., 2024a), and cross-problem generalization techniques (Liu et al., 2024; Berto et al., 2024b). Alternative solution strategies have also emerged, such as learning heuristics for ant colony optimization (Ye et al., 2024a; Kim et al., 2024), divide-and-conquer approaches (Li et al., 2021; Ye et al., 2024b; Zheng et al., 2024) and improvement methods (Kong et al., 2024; Hottung et al., 2025b; Ma et al., 2024).
However, significant challenges remain in bridging the gap between synthetic and real-world VRPs. Current approaches often struggle with asymmetric travel times, rich edge features, and complex constraints typical in practical scenarios. While some works have attempted to address these challenges by (partially) encoding edge information (Duan et al., 2020; Lei et al., 2022; Kwon et al., 2021; Drakulic et al., 2024b, a) most are still designed for and evaluated on simplified VRP variants with distances obtained in Euclidean space, with no distinction between travel distances and travel times, and with inefficient learning of neural representations. The development of NCO architectures capable of effectively processing and utilizing structural information from real-world routing scenarios remains an open challenge, which the proposed RRNCO aims to address.
3. Preliminaries
3.1. Vehicle Routing Problems
Vehicle Routing Problems (VRPs) are a class of combinatorial optimization problems that aim to find optimal routes for a fleet of vehicles serving a set of customers by minimizing a cost function. The simplest variant, the Traveling Salesman Problem (TSP), involves finding a minimum-cost Hamiltonian cycle in a complete graph $G=(V,E)$ with $n=|V|$ locations. Real-world applications typically extend this basic formulation with operational constraints such as vehicle capacity limits (CVRP) or time windows for service (VRPTW) (Vidal et al., 2014).
@@ 3.1. Vehicle Routing Problems
These problems are characterized by their input structure, consisting of node and edge features. Node features typically include location coordinates and customer demands, while edge features capture the relationships between locations. In real-world settings, these relationships are represented by distance and duration matrices, $D,T\in\mathbb{R}^{n\times n}$, where $d_{ij}$ and $t_{ij}$ denote the distance and travel time from location $i$ to $j$, respectively. Previous NCO approaches typically rely on Euclidean distances computed directly from location coordinates, avoiding the use of distance matrices entirely. While this simplification works for synthetic problems with symmetric distances ($d_{ij}=d_{ji}$, $t_{ij}=t_{ji}$), real-world instances are inherently asymmetric due to factors such as traffic patterns and road network constraints. This asymmetry, combined with the problem’s combinatorial nature, presents unique challenges for learning-based approaches, particularly in effectively encoding and processing the rich structural information present in edge features.
3.2. Solving VRPs with Generative Models
@@
VRPs can be solved as a sequential decision process, and deep generative models can be learned to efficiently do so (Wu et al., 2024b). Given a problem instance $\bm{x}$ containing both node features (such as coordinates, demands, and time windows) and edge features (distance and duration matrices $D,T$), we construct solutions through autoregressive generation. Our model iteratively selects the next location to visit based on the current partial route until all locations are covered. This construction process naturally aligns with how routes are executed in practice and allows the model to maintain feasibility constraints throughout generation.
In this work, we consider the encoder-decoder framework as in Kool et al. (2019). Formally, let $\theta=\{\theta_{f},\theta_{g}\}$ denote the combined parameters of our encoder and decoder networks. We learn a policy $\pi_{\theta}$ that maps input instances to solutions through:
(1a) $\bm{h}$ $=f_{\theta_{f}}(\bm{x}),$
(1b) $\pi_{\theta}(\bm{a}|\bm{x})$ $=\prod_{t=1}^{T}g_{\theta_{g}}(a_{t}|a_{t-1},\dots,a_{1},\bm{h}),$
@@
3.3. Training via Reinforcement Learning
@@
(2) $\max_{\theta}J(\theta)=\mathbb{E}_{\bm{x}\sim\mathcal{D}}\mathbb{E}_{\bm{a}\sim\pi_{\theta}(\cdot|\bm{x})}[R(\bm{a},\bm{x})],$
@@
4. Real-World VRP Dataset
Existing methodologies often require integrating massive raw datasets (e.g., traffic simulators and multi-source spatial data)(Barauskas et al., 2023) or rely on simplistic synthetic benchmarks, which are either resource-intensive or lack real-world complexity(Gunawan et al., 2021).
To address these limitations, we design a three-step pipeline to create a diverse and realistic vehicle routing dataset aimed at training and testing NCO models. First, we select cities worldwide based on multi-dimensional urban descriptors (morphology, traffic flow regimes, land-use mix) (Section 4.1). Second, we develop a framework using the Open Source Routing Machine (OSRM) (Luxen and Vetter, 2011) to create city maps with topological data, generating both precise location coordinates and their corresponding distance and duration matrices between each other (Section 4.2). Finally, we efficiently subsample these topologies to generate diverse VRP instances by adding routing-specific features such as demands and time windows, thus preserving the inherent spatial relationships while enabling the rapid generation of instances with varying operational constraints, leveraging the precomputed distance/duration matrices from the base maps (Section 4.3). The whole pipeline is illustrated in Fig. 2.
Figure 2. Overview of our RRNCO real-world data generation and sampling framework. We generate a dataset of real-world cities with coordinates and respective distance and duration matrices obtained via OSRM. Then, we efficiently subsample instances as a set of coordinates and their matrices from the city map dataset with additional generated VRP features.
4.1. City Map Selection
We select a list of 100 cities distributed across six continents, with 25 in Asia, 21 in Europe, 15 each in North America and South America, 14 in Africa, and 10 in Oceania. The selection emphasizes urban diversity through multiple dimensions, including population scale (50 large cities ¿1M inhabitants, 30 medium cities 100K-1M, and 20 small cities ¡100K), infrastructure development stages, and urban planning approaches. Cities feature various layouts, from grid-based systems like Manhattan to radial patterns like Paris and organic developments like Fez, representing different geographic and climatic contexts from coastal to mountain locations.
We prioritized cities with reliable data availability while balancing between globally recognized metropolitan areas and lesser-known urban centers, providing a comprehensive foundation for evaluating vehicle routing algorithms under diverse real-world conditions. Moreover, by including cities from developing regions, we aim to advance transportation optimization research that could benefit underprivileged areas and contribute to their socioeconomic development (see Appendix A for more details).
4.2. Topological Data Generation Framework
Geographic boundary information
(3) $d=2R\cdot\arcsin\left(\sqrt{\sin^{2}\left(\frac{\Delta\phi}{2}\right)+\cos(\phi_{1})\cos(\phi_{2})\sin^{2}\left(\frac{\Delta\lambda}{2}\right)}\right)$
Point sampling from road networks
Figure 3. Our proposed RRNCO model for real-world routing.
Travel information computation
The travel information computation component leverages a locally hosted Open Source Routing Machine (OSRM) server (Luxen and Vetter, 2011) to calculate real travel distances and durations between sampled points, ensuring full reproducibility of results. Through the efficient get_table function in our router implementation via the OSRM table service [^2] https://project-osrm.org/docs/v5.24.0/api/#table-service, we can process a complete 1000x1000 origin-destination matrix within 18 seconds, making it highly scalable for urban-scale analyses. In contrast to commercial API-based approaches that require more than 20 seconds for 350×350 matrices (Ali and Saleem, 2024), our open-source local OSRM implementation achieves the same computations in approximately 5 seconds. Additionally, it enables the rapid generation of multiple instances from small datasets with negligible computational cost per iteration epoch.
The RRNCO framework finally processes this routing data through a normalization strategy that addresses both unreachable destinations and abnormal travel times. This step captures real-world routing complexities, including one-way streets, turn restrictions, and varying road conditions, resulting in asymmetric distance and duration matrices that reflect actual urban travel patterns. All computations are performed locally [^3] Our framework can also be extended to include real-time commercial map API integrations and powerful traffic forecasting to obtain better-informed routing (Wang and Xu, 2011; Qu and Wu, 2025), which we leave as future works., allowing for consistent results and independent verification of the analysis pipeline.
4.3. VRP Instance Subsampling
From the large-scale city base maps, we generate diverse VRP instances by subsampling a set of locations along with their corresponding distance and duration matrices, allowing us to generate an effectively unlimited number of instances while preserving the underlying structure.
The subsampling process follows another three-step procedure:
1. Index Selection: Given a city dataset containing $N_{\text{tot}}$ locations, we define a subset size $N_{\text{sub}}$ representing the number of locations to be sampled for the VRP instance. We generate an index vector $\mathbf{s}=(s_{1},s_{2},\dots,s_{N_{\text{sub}}})$ where each $s_{i}$ is drawn from $\{1,\dots,N_{\text{tot}}\}$, ensuring unique selections.
2. Matrix Subsampling: Using $\mathbf{s}$, we extract submatrices from the precomputed distance matrix $D\in\mathbb{R}^{N_{\text{tot}}\times N_{\text{tot}}}$ and duration matrix $T\in\mathbb{R}^{N_{\text{tot}}\times N_{\text{tot}}}$, forming instance-specific matrices $D_{\text{sub}}=D[\mathbf{s},\mathbf{s}]\in\mathbb{R}^{N_{\text{sub}}\times N_{\text{sub}}}$ and $T_{\text{sub}}=T[\mathbf{s},\mathbf{s}]\in\mathbb{R}^{N_{\text{sub}}\times N_{\text{sub}}}$, preserving spatial relationships among selected locations.
3. Feature Generation: Each VRP can have different features. For example, in Asymmetric Capacitated VRP (ACVRP) we can generate a demand vector $\mathbf{d}\in\mathbb{R}^{N_{\text{sub}}\times 1}$, such that $\mathbf{d}=(d_{1},d_{2},\dots,d_{N_{\text{sub}}})^{\top}$, where each $d_{i}$ represents the demand at location $s_{i}$. Similarly, we can extend to ACVRPTW (time windows) represented as $\mathbf{W}\in\mathbb{R}^{N_{\text{sub}}\times 2}$, where $\mathbf{W}=\{(w_{1}^{\text{start}},w_{1}^{\text{end}}),\dots,(w_{N_{\text{sub}}}^{\text{start}},w_{N_{\text{sub}}}^{\text{end}})\}$, defining the valid service interval for each node.
Unlike previous methods that generate static datasets offline (Duan et al., 2020; Ali and Saleem, 2024), our RRNCO generation framework dynamically generates instances on the fly in few milliseconds, reducing disk memory consumption while maintaining high diversity.
Fig. 2 illustrates the overall process, showing how a city map is subsampled using an index vector to create VRP instances with distance and duration matrices enriched with node-specific features such as demands and time windows. Our approach allows us to generate a (arbitrarily) large number of problem instances from a relatively small set of base topology maps totaling around $1.5$GB, in contrast to previous works that required hundreds of gigabytes of data to produce just a few thousand instances.
5. Real-World Routing Model
Our RRNCO model addresses the unique challenges of real-world routing, where conventional methods struggle with asymmetric routing attributes such as travel times and distances. To overcome these issues, we introduce two key innovations: (1) a spatial feature integration mechanism that fuses coordinate and distance matrix information through a contextual gating system, and (2) Neural Adaptive Biases (NAB), replacing heuristic adaptation rules within the Adaptation Attention Free Module (AAFM) (Zhou et al., 2024b) to better capture complex relationships.
The model employs an encoder-decoder architecture (see Eq. 1), where the encoder constructs comprehensive node representations integrating local and global routing contexts, and the decoder sequentially generates solutions. Our main contributions focus on improving the encoder’s ability to process real-world routing complexities while maintaining computational efficiency.
5.1. Encoder
5.1.1. Initial Embedding Module

### v2

# RRNCO: Towards Real-World Routing with Neural Combinatorial Optimization

@@
Abstract
The practical deployment of Neural Combinatorial Optimization (NCO) for Vehicle Routing Problems (VRPs) is hindered by a critical sim-to-real gap. This gap stems not only from training on oversimplified Euclidean data but also from node-based architectures incapable of handling the node-and-edge-based features with correlated asymmetric cost matrices, such as those for real-world distance and duration. We introduce RRNCO, a novel architecture specifically designed to address these complexities. RRNCO’s novelty lies in two key innovations. First, its Adaptive Node Embedding (ANE) efficiently fuses spatial coordinates with real-world distance features using a learned contextual gating mechanism. Second, its Neural Adaptive Bias (NAB) is the first mechanism to jointly model asymmetric distance, duration, and directional angles, enabling it to capture complex, realistic routing constraints. Moreover, we introduce a new VRP benchmark grounded in real-world data crucial for bridging this sim-to-real gap, featuring asymmetric distance and duration matrices from 100 diverse cities, enabling the training and validation of NCO solvers on tasks that are more representative of practical settings. Experiments demonstrate that RRNCO achieves state-of-the-art performance on this benchmark, significantly advancing the practical applicability of neural solvers for real-world logistics. Our code, dataset, and pretrained models are available at https://github.com/ai4co/real-routing-nco.
1 Introduction
@@
Vehicle routing problems (VRPs) are combinatorial optimization (CO) problems that represent foundational challenges in logistics and supply chain management, directly impacting operations across diverse sectors, including last-mile delivery services, disaster response management, and urban mobility. These NP-hard optimization problems require determining optimal routes for a fleet of vehicles while satisfying various operational constraints. While several traditional methods have been developed over decades (Laporte and Nobert, 1987; Vidal, 2022; Wouda et al., 2024; Perron and Furnon, 2023; Applegate et al., 2003; Wouda and Lan, 2023), these often face challenges in real-world applications. Their computational complexity makes them impractical for large-scale and real-time applications. Moreover, they often require careful parameter tuning, problem-specific adaptations, significant domain expertise, and lengthy development. With the global logistics market exceeding the $10$ trillion USD mark in 2025 (Research and Markets, 2024), improvements in routing efficiency can yield substantial cost savings and environmental benefits.
@@
Neural Combinatorial Optimization (NCO) has emerged as a promising paradigm for solving CO problems such as the VRP (Bengio et al., 2021; Wu et al., 2024a). By automatically learning heuristics directly from data, i.e., by using Reinforcement Learning (RL), NCO approaches for VRPs can potentially overcome the limitations of traditional methods by providing efficient solutions without requiring extensive domain expertise and by providing more scalable solutions (Kool et al., 2019; Zhou et al., 2023). Recent advances in NCO have demonstrated impressive results on synthetic VRP instances, suggesting the potential for learning-based approaches to achieve significant impact in real-world logistics optimization (Kwon et al., 2020; Luo et al., 2024; Ye et al., 2024b; Hottung et al., 2025).
However, while real-world VRPs encompass various dynamic and operational complexities, the transition from synthetic to practical applications faces a primary topological challenge.
@@
Firstly, most existing NCO research primarily relies on simplified synthetic datasets for both training and testing that fail to capture the foundational asymmetries of real-world road networks, particularly asymmetric travel times and distances arising from road networks with diverse conditions (Osaba, 2020; Thyssens et al., 2023). Hence, a comprehensive framework for real-world data generation is needed to bridge this gap. Secondly, most current NCO architectures are based on the node-based transformer paradigm (Vaswani et al., 2017) and, as such, are not designed to effectively and efficiently embed the rich edge features and structural information present in real-world routing problems, limiting their practical applicability (Kwon et al., 2021). A new neural approach capable of effectively encoding information such as asymmetric durations and distances is thus needed to bridge the sim-to-real gap.
Figure 1: [Left] Most NCO works consider simplified Euclidean settings. [Right] Our work models real-world instances where durations and travel times can be asymmetric.
Our Real Routing NCO (RRNCO) bridges the critical topological gap between simplified NCO research and real-world routing applications —as illustrated in Fig. 1—through architectural innovations that specifically target the asymmetric nature of practical routing problems while serving as an extensible foundation for complex environments. We make two fundamental contributions. First, on the modeling side, we introduce a novel neural architecture with two key technical innovations: (i) an Adaptive Node Embedding (ANE) that dynamically fuses coordinates and distance information via learned contextual gating and probability-weighted sampling; and (ii) a Neural Adaptive Bias (NAB), the first mechanism to jointly model asymmetric distance and duration matrices within a deep routing framework, guiding our Adaptation Attention Free Module (AAFM). To validate our approach, we construct a comprehensive benchmark dataset from 100 diverse cities, featuring real-world asymmetric distance and duration matrices from OpenStreetMap (OpenStreetMap contributors, 2025).
Our contributions: (1) A novel NCO architecture (RRNCO) with ANE and NAB to natively handle real-world routing asymmetries. (2) An extensive, open-source VRP dataset from 100 cities with asymmetric matrices. (3) State-of-the-art empirical results on realistic VRP instances. (4) Open-source code and data to foster reproducible research.
2 Related Works
Neural Combinatorial Optimization (NCO)
Neural approaches to combinatorial optimization learn heuristics directly from data, reducing reliance on domain expertise (Bengio et al., 2021). Methods are typically classified as construction or improvement. Construction methods sequentially generate solutions, pioneered by Pointer Networks (Vinyals et al., 2015) and now led by Transformer-based autoregressive models (Kool et al., 2019; Kwon et al., 2020) for their strong ability to capture complex structures. Non-autoregressive variants predict edge-probability heatmaps in a single pass (Joshi et al., 2020), with later work enhancing performance via stronger models and search strategies (Ye et al., 2024a; Sun and Yang, 2024; Xia et al., 2024; Kim et al., 2025). Improvement methods iteratively refine an initial solution through learned operators or policies (Hottung and Tierney, 2019; Ma et al., 2023; Hottung et al., 2021; Son et al., 2023; Li et al., 2023; Chalumeau et al., 2023; Kim et al., 2021; Ma et al., 2021; Ye et al., 2024b; Zheng et al., 2024). Our work targets autoregressive construction, striking an effective balance between speed and solution quality for real-world logistics.
Vehicle Routing Problem (VRP) Datasets
A significant gap exists between NCO research and real-world applicability, largely due to the datasets used for training and evaluation. For decades, the community has relied on established benchmarks like TSPLIB (Reinelt, 1991) and CVRPLIB (Lima et al., 2014). While invaluable for standardization, these datasets are typically based on symmetric Euclidean distances, assuming travel costs are equal in both directions ($d_{ij}=d_{ji}$). This simplification fails to capture the inherent asymmetry of real road networks caused by one-way streets, traffic patterns, and turn restrictions (Osaba, 2020). Some recent works have attempted to create more realistic datasets (Duan et al., 2020; Ali and Saleem, 2024), but they suffer from critical limitations for NCO research: they often rely on proprietary, commercial APIs, are static and cannot be generated online (a key requirement for data-hungry RL agents), can be slow to generate, and are not always publicly released. Furthermore, they often omit crucial information like travel durations, which can be decoupled from distance in real traffic. Our work directly addresses these gaps by providing a fast, open-source, and scalable data generation framework that produces asymmetric distance and duration matrices from real-world city topologies.
NCO for VRPs
The application of NCO to VRPs has evolved from early adaptations of recurrent models (Nazari et al., 2018) to the now-dominant Transformer-based encoder-decoder architectures (Kool et al., 2019; Kwon et al., 2020; Kim et al., 2022; Luo et al., 2023; Zhou et al., 2024b; Huang et al., 2025; Luo et al., 2025; Berto et al., 2025b). These models have demonstrated impressive performance but are fundamentally node-centric; their attention mechanisms operate on node embeddings, making it non-trivial to incorporate rich structural information contained in edge features like a full distance matrix. This limitation is a primary contributor to the sim-to-real gap. To address this, some works have explored encoding edge information. GCN-based approaches (Duan et al., 2020) and attention via row and column embeddings of MatNet (Kwon et al., 2021) introduced early ways to handle asymmetry, with GOAL (Drakulic et al., 2024) incorporating edge data with cross-attention. While promising, existing methods typically handle only a single cost matrix (e.g., distance) and fail to leverage the correlated modalities of real-world routing (distance, duration, geometry). Efficiently fusing multiple asymmetric edge features remains an open challenge. Our model, RRNCO, addresses this with Adaptive Node Embeddings (ANE) and a Neural Adaptive Bias (NAB) mechanism that learns a unified routing context from distance, duration, and angle.
3 Preliminaries
3.1 Vehicle Routing Problems
@@
Vehicle Routing Problems (VRPs) are a class of combinatorial optimization problems that aim to find optimal routes for a fleet of vehicles serving a set of customers by minimizing a cost function. The simplest variant, the Traveling Salesman Problem (TSP), involves finding a minimum-cost Hamiltonian cycle in a complete graph $G=(V,E)$ with $n=|V|$ locations. Real-world applications typically extend this basic formulation with operational constraints such as vehicle capacity limits (CVRP) or time windows for service (VRPTW) (Vidal et al., 2014). These problems are characterized by their input structure, consisting of node and edge features. Node features typically include location coordinates and customer demands, while edge features capture the relationships between locations. In real-world settings, these relationships are represented by distance and duration matrices, $D,T\in\mathbb{R}^{n\times n}$, where $d_{ij}$ and $t_{ij}$ denote the distance and travel time from location $i$ to $j$, respectively. Previous NCO approaches typically rely on Euclidean distances computed directly from location coordinates, avoiding the use of distance matrices entirely. While this simplification works for synthetic problems with symmetric distances ($d_{ij}=d_{ji}$, $t_{ij}=t_{ji}$), real-world instances are inherently asymmetric due to factors such as traffic patterns and road network constraints. This asymmetry, combined with the problem’s combinatorial nature, presents unique challenges for learning-based approaches, particularly in effectively encoding and processing the rich structural information present in edge features.
3.2 Solving VRPs with Generative Models
@@
VRPs can be solved as a sequential decision process, and deep generative models can be learned to efficiently do so (Wu et al., 2024b). Given a problem instance $\bm{x}$ containing both node features (such as coordinates, demands, and time windows) and edge features (distance and duration matrices $D,T$), we construct solutions through autoregressive generation. Our model iteratively selects the next location to visit based on the current partial route until all locations are covered. This construction process naturally aligns with how routes are executed in practice and allows the model to maintain feasibility constraints throughout generation. In this work, we consider the encoder-decoder framework as in Kool et al. (2019). Formally, let $\theta=\{\theta_{f},\theta_{g}\}$ denote the combined parameters of our encoder and decoder networks. We learn a policy $\pi_{\theta}$ that maps input instances to solutions through:
$\bm{h}$ $=f_{\theta_{f}}(\bm{x}),$ (1a)
$\pi_{\theta}(\bm{a}|\bm{x})$ $=\prod_{t=1}^{T}g_{\theta_{g}}(a_{t}|a_{t-1},\dots,a_{1},\bm{h}),$ (1b)
@@
3.3 Training via Reinforcement Learning
@@
$\max_{\theta}J(\theta)=\mathbb{E}_{\bm{x}\sim\mathcal{D}}\mathbb{E}_{\bm{a}\sim\pi_{\theta}(\cdot|\bm{x})}[R(\bm{a},\bm{x})],$ (2)
@@
4 Real-World Routing Model
Figure 2: Our proposed RRNCO model for real-world routing.
Our model addresses real-world routing challenges where conventional methods struggle with asymmetric attributes like travel times and distances. We introduce two key innovations: (1) Adaptive Node Embedding with probability-weighted distance sampling - efficiently integrating spatial coordinates with asymmetric distances through learned contextual gating, avoiding full distance matrix processing while preserving asymmetric relationships, and (2) Neural Adaptive Bias (NAB) - the first learnable mechanism to jointly model asymmetric distance and duration matrices in deep routing architectures, replacing hand-crafted heuristics in AAFM (Zhou et al., 2024a) with data-driven contextual biases. The model uses an encoder-decoder architecture where the encoder builds comprehensive node representations and the decoder generates solutions sequentially, with our contributions focusing on enhancing the encoder’s real-world routing capabilities while maintaining efficiency.
4.1 Encoder
4.1.1 Adaptive Node Embedding